| Sketch the curve of the function f(x) = + unction f(x) = * [r(x) = 2*, S(x) = 205*] Do this by determining the following
Answer #1

4 domain XF ( 9) UC 0,-) - y-interlalba + When y=0 - x=-1 and x=1 y.inter opt - DNE - Vertica assy ontstoke (1907 - HiA - lys